OpenStudy (youngstudier):

Laketown had 215 inches of snowfall last year, but only 155 inches this year. What was the percent decrease in snowfall for Laketown to the nearest percent? A. 28% B. 39% C. 72% D. 139% Not sure if it's C or D.

OpenStudy (youngstudier):

@texaschic101

OpenStudy (youngstudier):

@unkownbeast911

OpenStudy (youngstudier):

Help please.

OpenStudy (youngstudier):

@iGreen

OpenStudy (youngstudier):

@Mahmood786123

OpenStudy (mahmood786123):

gimme a seccond

OpenStudy (mahmood786123):

do 215 minus 155

OpenStudy (youngstudier):

60

OpenStudy (mahmood786123):

215-155=60

OpenStudy (mahmood786123):

good now change 60 to a percentage

OpenStudy (mahmood786123):

D

OpenStudy (mahmood786123):

I think

OpenStudy (mahmood786123):

Hope I helped :)

OpenStudy (youngstudier):

oh LOL, I did 215/155 = x/100 155x=21,500 x=138.70967 x=139

OpenStudy (youngstudier):

Thanks